Two planes start from the same point and fly in opposite directions. The first plane is flying 50 mph slower than the second plane. In 3.5 h, the planes are 2100 mi apart. Find the rate of each plane.
Found 2 solutions by stanbon, ikleyn:
Answer by stanbon(75887)

Two planes start from the same point and fly in opposite directions. The first plane is flying 50 mph slower than the second plane. In 3.5 h, the planes are 2100 mi apart. Find the rate of each plane.
-----------------------
faster plane DATA:
time = 3.5h ; rate = x mph ; distance = 3.5x miles
slower plane DATA:
time = 3.5h ; rate = x-50 mph ; distance = 3.5(x-50) miles
----------------------
Equation:
3.5x + 3.5x - 3.5*50 = 2100
7x = 2275
x = 325 mph (faster plane speed)
x-50 = 275 mph (slower plane speed)
-------------
Cheers,
Stan H.
-----------

Answer by ikleyn(52832)

.
Two planes start from the same point and fly in opposite directions. The first plane is flying 50 mph slower than the second plane.
In 3.5 h, the planes are 2100 mi apart. Find the rate of each plane.
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

3.5*u + 3.5*(u+50) = 2100.


This is the "distance" equation.

The unknown "u" is the average speed of the slower plane.

Solve for "u".

Happy calculations !!